OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −3

OpenStudy (chaise):

4x − 2y if x = −1 and y = −3 =4*-1 - 2x-3 =-4 - -6 =-4 + 6 =+2

OpenStudy (anonymous):

4(-1) - 2(-3) -4 + 6 = 2